OpenStudy (clamin):

PLEASE HELP!! MEDAL The function A (b) relates to the area of trapezoid with a given height of 10 and one base length of 7 with the length of its other base. It takes as input the other base value, and returns as output the area of the trapezoid. A (b)= 10 × b+7/ 2 Which equation below represents the inverse function B (a) which takes the trapezoid's area as input and returns as output the length of the other base?

OpenStudy (clamin):

A. B (a) = a/5 -7 B. B (a) = a/5 +7 C. B (a)= a/7 -5 D. B (a) = a/7 +5

OpenStudy (anonymous):

I believe \[A(b) = \frac{1}{2} \times 10 \times (b+7)\] The question asks you to find the inverse of the function which means "write b in terms of a" Multiply both sides by 2 and divide both sides by 10 \[\frac{2A}{10}=b+7\] Subtract both sides by 7 \[\frac{a}{5}-7=b(A)\]